OpenStudy (anonymous):

(someone please help me D: I'll give a medal and I will fan you ) 3. What is the slope of the line that passes through the pair of points? (1, 7), (10, 1)

OpenStudy (anonymous):

\[a. \frac{ 2 }{ 3} \] b. \[\frac{ -2 }{ 3}\] c. \[\frac{ -3 }{ 2 }\] d. \[\frac{ 2 }{ 3}\]

OpenStudy (praxer):

let (x,y) and (x',y') be two points and it is joined by a line than the slope to the line is given by the formula slope= (y'-y)/(x'-x) http://www.mathwarehouse.com/algebra/linear_equation/images/slope-of-a-line2.gif Once you get this. You will be able to do it smoothly.. Hope it helps. :)

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

slope is the rate of change change in y / corresponding change in x for 2 given points (x1, y1) and (x2, y2) slope = y2 -y1 /x2 - x1 for this problem (1, 7), (10, 1) make the substitution and solve

OpenStudy (triciaal):

(1-7 )/ (10-1) = -6/9 reduce to lowest term so divide by 3/3 to get -2/3

OpenStudy (anonymous):

*writes it all down*

OpenStudy (anonymous):

thank you :)

OpenStudy (triciaal):

b

OpenStudy (triciaal):

*writes it all down* what does this mean?

OpenStudy (anonymous):

it means that I wrote all what you said so I can remember. :)

OpenStudy (triciaal):

OK slope is the relationship of how the dependent variable changes with the corresponding change in the independent variable. This is written as a ratio of integers.

OpenStudy (anonymous):

oooo oki thanks :)

OpenStudy (triciaal):

you are welcome